OpenStudy (monkey*):

Precal Help! Solve algebraically. Y+2x=7 Y=5-x

OpenStudy (anonymous):

the second equation is already solved for y, so substitute it into the first and you get \[5-x+2x=7\] solve for x, then go back to the 2nd equation to solve for y

OpenStudy (monkey*):

X=2

OpenStudy (anonymous):

right, now put the 2 into the second equation for x and solve for y

OpenStudy (monkey*):

5-2+2(2)=7 like this?

OpenStudy (anonymous):

not quite. You want to use the original equations for this part. If you use the first one. y + 2x = 7 Now sub the 2 for x y + 2(2) = 7 Solve for y If you use the second. y = 5 - x sub the 2 for x y = 5 - 2 You get the same value for y with either equation

OpenStudy (monkey*):

So the solutions would be (2,3)?

OpenStudy (anonymous):

yes

OpenStudy (monkey*):

Okay, thank you so much!